### v1

# OmniACT: A Dataset and Benchmark for Enabling Multimodal Generalist Autonomous Agents for Desktop and Web

Abstract
footnotetext: $\dagger$ These authors contributed equally. The order is determined by dice rolling.
For decades, human-computer interaction has fundamentally been manual. Even today, almost all productive work done on the computer necessitates human input at every step. Autonomous virtual agents represent an exciting step in automating many of these menial tasks. Virtual agents would empower users with limited technical proficiency to harness the full possibilities of computer systems. They could also enable the efficient streamlining of numerous computer tasks, ranging from calendar management to complex travel bookings, with minimal human intervention. In this paper, we introduce OmniACT, the first-of-a-kind dataset and benchmark for assessing an agent’s capability to generate executable programs to accomplish computer tasks. Our scope extends beyond traditional web automation, covering a diverse range of desktop applications. The dataset consists of fundamental tasks such as “Play the next song”, as well as longer horizon tasks such as “Send an email to John Doe mentioning the time and place to meet”. Specifically, given a pair of screen image and a visually-grounded natural language task, the goal is to generate a script capable of fully executing the task. We run several strong baseline language model agents on our benchmark. The strongest baseline, GPT-4, performs the best on our benchmark However, its performance level still reaches only 15% of the human proficiency in generating executable scripts capable of completing the task, demonstrating the challenge of our task for conventional web agents. Our benchmark provides a platform to measure and evaluate the progress of language model agents in automating computer tasks and motivates future work towards building multimodal models that bridge large language models and the visual grounding of computer screens.
Figure 1: OmniACT dataset and benchmark for enabling autonomous human-computer interaction agents. The left shows an image paired with a natural language task description as the input, the right shows the resulting action script to be executed on the screen. Examples are presented from Stocks, Apartments.com, and Weather application.
1 Introduction
Performing computer tasks based on natural language instructions has been a long-standing goal of artificial intelligence [45]. One concrete objective in the line of research is to develop generalist agents that can assist humans in doing computer tasks [19], such as “Order a pizza from Domino’s” or “Write a message to John.” The agent should be able to open the application and perform the task. Executing these actions on a personal computer involves a sequence of interactions with a mouse and keyboard. For example, the simple task of writing an email involves hovering over the application icon, clicking it, clicking the ‘New Email’ button, writing the content of the email, and clicking send. Successfully sending an email requires accurately predicting the correct action at each step and accurately executing it, which is a herculean task even for the best agents today [13].
A generalist agent for computer tasks must understand natural language instructions, process visual screenshots, and produce the correct sequence of actions to be performed to achieve the intended task. Several existing approaches focus on building agents based on the HTML model [37, 9, 57]. However, this approach introduces several challenges and constraints. These agents are limited to web applications and often struggle with complex or long-context HTML code. They cannot interact with native desktop applications or perform tasks that span multiple applications, like drafting an email using text from a code editor, without significant alterations. Furthermore, HTML-based agents, which are inherently powered by text-only language models, typically underperform in tasks requiring visual cues, such as identifying and clicking a blue button on a desktop’s top-right corner. In contrast, humans can easily understand UI elements like dropdown menus, typable areas, redirections, and options with just a glance.
Towards the goal of developing a generalist autonomous agent with robust visual and user interface (UI) understanding capabilities, we introduce a new task and dataset, OmniACT, containing over 9.8K pairs of images and instructions (Figure 1) across different operating systems and the web. This dataset includes screenshots of various UI screens and corresponding natural language instructions. The objective of these instructions is to generate executable commands using the PyAutoGUI Python library [1]. PyAutoGUI enables the automation of the mouse and keyboard operations, which helps to facilitate interactions with various native applications across macOS, Windows, and Linux. This simplifies completing specified tasks across different web domains and native desktop applications.
We evaluate several language model-based agent baselines on this dataset, including LLaMA [43], Vicuna [7], Palmyra-X (43B) [2], InstructPalmyra-30B [41], GPT 3.5, and GPT-4 [30]. We experiment with fine-tuning Vicuna-13B and LLaMA-13B models using QLoRA [10]. We also benchmark multimodal baseline LLaVa-v1.5-7B, LLaVa-v1.5-13B [43] and GPT-4-vision-preview [50] for the task. Our findings highlight the necessity for a multimodal model capable of executing these tasks, and our analysis provides insights into promising future work in the space. Our key contributions are outlined as follows:
  1. 1.
    We release a novel dataset of desktop and website applications consisting of over 9.8K natural language tasks, UI screens, and corresponding code snippets collected through human annotation. We introduce custom performance metrics tailored for computer tasks.
  2. 2.
    We propose DetACT, a module for creating textual representations of the screen using signals from OCR, color, and icon-template matching.
  3. 3.
    We conduct a comprehensive benchmark and analysis of state-of-the-art LLMs and multimodal models on our benchmark. Our results show that OmniACT is a challenging task for even the best LLM agents today, and existing models are far below human performance.
2 Related Work
Datasets Size Env Type
Task
Heterogeneity
Real-World
Portayal
Executional
Correctness
Supports
Desktop
Apps
Continuous Scale
Adaptive
Evaluation
Task
WebArena [57] 812 Web Yes Yes Yes No No Web Navigation
Mind2Web [9] 2350 Web Yes Yes No No No Web Navigation
WebShop [51] 12000 Products Web No No Yes No No Web Navigation
RUSS [49] 80 Web Yes Yes No No No Web Navigation
WebSRC [6] 2735 Web Yes Yes - No No QA
MiniWoB++ [16] 100
Mobile
Websites
No No Yes No No Web Navigation
PixelHelp [21] 187 Mobile Yes Yes No No No UI Grounding
MetaGUI  [39] 1125 Mobile Yes Yes Yes No No Mobile Navigation
MoTIF [5] 756 Mobile Yes Yes Yes No No Mobile Navigation
AITW [32] 715142 Mobile and Web Yes Yes Yes No No
Mobile/Web
Navigation
OmniACT (Ours) 9802 Desktop and Web Yes Yes Yes Yes Yes Code Generation
Table 1: Comparison of OmniACT with other related benchmarks.
2.1 UI Understanding
User interface (UI) understanding has garnered interest from researchers in the machine learning and human-computer interaction communities, evolving with various models focusing on understanding the semantics of mobile and web user interfaces. UIBert [3], PixelBERT [15], ActionBert [14], VUT [23], Screen2Words [44], WidgetCaptioning [22] and Pix2Act [36] are notable models in this area. They propose approaches for learning the user-interface semantics of the mobile screen using the image and view hierarchy. These models have demonstrated effectiveness in tasks like capability prediction, screen segmentation and understanding, and screen caption generation. Lexi [4] and Spotlight [20] propose models that use vision-only inputs to minimize the reliance on metadata such as view hierarchy. Furata et al. [11] demonstrates the use of fine-tuning for multimodal web navigation. The majority of machine learning models trained for UI understanding leverage the Rico dataset [8] and its extensions, which contain 64,462 unique Android screens and metadata. In addition, Banerjee et al. [4] released the UICaptions dataset, which consists of diverse image-captions pairs across a wide range of applications. PixelHelp [21] also released a corpus to train models that can interpret natural language instructions and map them to mobile user interface actions.
2.2 Autonomous Computer Agents
The advent of large language models (LLMs) has been pivotal in the rapid advancement of agents that operate on web pages. Recent research such as ViperGPT [40] Chameleon [26], RCI Agent [17], VisProg [12], and [28] employ LLMs for planning or action prediction in developing autonomous agents. Benchmark datasets, such as MiniWoB [37], WebShop [51], Macaw-LLM [27], ASH-Prompting [38] Mind2Web [9] and WebArena [57] have also been proposed to measure the ability of LLM-based agents in automating web tasks. These methods mainly involve agents that operate on a text-based Document Object Model (DOM) of HTML scripts. This limits their understanding of screen context, which is crucial for the model’s decision-making and action-taking processes. To address this limitation, Rawles et al. [32] released Android in the Wild, a dataset comprising screens, natural language instructions, and corresponding actions. Following this,  [54] proposed a multimodal model, AutoUI, which is designed to build an agent on the Android in the Wild dataset confined to the Android ecosystem.
Current benchmarks for autonomous agents focus mainly on the Web or Android environments, posing challenges for tasks involving desktop applications or spanning multiple applications beyond the web domain. The absence of established benchmarks and datasets in this area, coupled with basic methods for extracting user interface (UI) elements, underscores the need for significant progress in developing more versatile autonomous agents capable of handling diverse tasks beyond the current scope. To highlight the unique features that OmniACT introduces in the assessment of capable autonomous agents, we provide a comparison between the existing benchmarks and our proposed benchmark, OmniACT, in Table 1.
Figure 2: Data Collection Pipeline. (1) We select over 60 applications and websites to ensure diversity, (2) segment the screen through human-annotated bounding boxes, (3) label the bounding boxes based on functionality, (4) ask student volunteers to come up with tasks, given a screen image, and (5) reverse map the textual labels to coordinates and filter the scripts based on execution and syntax.
3 OmniACT
We introduce a novel dataset and benchmark, OmniACT, which measures the performance of autonomous agents on both web and desktop applications. Compared to previous benchmarks which focus on text-based reasoning [37, 57, 9, 51, 16], our benchmark aims to measure multimodal agents that bridge large language model planners and UI understanding vision models. OmniACT can be accomplished as a standalone task as it is not under a mock environment.
All actions that a human can execute on the computer can be encoded in the PyAutoGUI [1] Python framework. This framework allows a user to execute keyboard and mouse operations by running Python code. The PyAutoGUI code to execute these tasks is shown in the third column of Figure 1. For other computer tasks, the PyAutoGUI library provides functions such as ‘press’, ‘write’, and ‘scroll’ which can be used to execute the task. Our dataset consists of parallel data of natural language tasks, UI screenshots, and ground truth PyAutoGUI scripts that achieve successful execution.
3.1 Task Formulation
[…]
Figure 1 illustrates sample tasks and corresponding outputs for three applications within OmniACT: (1) Stocks (MacOS), (2) Apartments.com (web page), and (3) Weather (MacOS). The first column depicts the input image, and the second column shows the natural language task that is to be executed on the current screen. To execute these tasks, a user must accurately perform a series of operations using the mouse and keyboard. For example, to check the rate of change in Google’s stock price over the last month, the mouse has to be moved to the last month and dragged while holding the left-click button to the current month.
3.2 Dataset Preparation
To prepare our dataset, we followed a pipelined approach, as summarized in Figure 2. We first selected a variety of applications and websites. For each application or website, we created bounding boxes around key UI elements and labeled them according to their functionality, which is crucial for assisting human annotators in writing accurate PyAutoGUI scripts. After each script is written, we converted the labels back into numeric coordinates, allowing us to align the scripts precisely with the locations of the UI elements. Finally, we thoroughly reviewed each script, focusing on its executability and adherence to syntax standards. This ensured the high quality and functionality of our dataset, making it a valuable resource for training and evaluating autonomous agents.
Type Action %
Click 63.73
Double Click 0.58
Right Click 0.77
Move/Hover 1.85
Drag 0.29
Scroll 1.68
Mouse Horizontal Scroll 0.17
Press 16.28
Hotkey 3.00
Keyboard Write 11.65
Table 2: Action type supported by OmniACT and the number of instances for each of the actions in the dataset.
3.2.1 Application/Website Selection
To test the computer agents’ generalization ability across different tasks, we collect tasks across multiple domains on both desktop and web applications. In total, we collect and annotate 9802 data points (Table 3), with the split between desktop and web applications approximately 3:1. The emphasis on desktop applications, which do not contain Document Object Model (DOM) hierarchies unlike HTML-based web pages, presents a more complex multimodal challenge where visual cues are crucial. We collect tasks from applications within the three most popular operating systems. We select 22 native applications from MacOS, and 8 each from Linux and Windows. We annotate roughly 3 to 4 screens for every application. The full list of applications is provided in the Appendix.
Many common computer tasks today are still performed through web applications, so we also collect 3-4 screenshots from 27 different web applications. To ensure diversity in task intents, we categorize these tasks into one of the following 6 categories: (1) Shopping, (2) Entertainment, (3) Service, (4) Government, (5) Travel, (6) Health. Inspired by the methodology of [9], these categories were selected to cover a wide range of user intents and functionalities.
3.2.2 UI Screen Segmentation
To collect gold-standard data, we first annotate and segment the screen by identifying the bounding boxes present on the screen. We employ slightly different techniques for web and desktop applications to create the bounding boxes:
  1. 1.
    Desktop Applications: We build a custom annotation interface based on PyQt5111https://pypi.org/project/PyQt5/ to create bounding boxes manually over a screen image using a simple drag-and-click mechanism. This custom interface expedites the process and allows us to get highly accurate gold-label data points for desktop images.
  2. 2.
    Websites: For webpages, we write JavaScript code to extract all interactable (click, hover, type, etc.) regions from the HTML source code. We also extract banners, dropdowns, submit, and radio buttons from the screen. We filter the elements to retain only those that are visible and interactable within the screen.
3.2.3 Functionality Tagging
To map each bounding box to its correct functional description, we leverage Amazon MTurk workers (see details in Appendix), who are given an image with a bounding box and are required to write the correct description or label of the bounding box’s function. For example, given an image of an Amazon webpage with a search bar, the annotator labels it as “find-product-search-bar”. The logical descriptions are used to create tasks in a structured manner without the need to identify individual bounding box coordinates.
3.2.4 Task Creation
Our approach for each screen involves utilizing all human-annotated bounding boxes and their labels to create tasks that can be executed within the confines of a single screen. These tasks are designed to be visually grounded in order to measure the capabilities of multimodal agents. We plan to release the bounding box and their corresponding labels as the metadata for evaluation purposes.
For dataset compilation, college students with basic Python programming skills served as annotators, accessing API references for PyAutoGUI and examples of potential tasks. Each student generated multiple tasks, each accompanied by three alternative natural language reformulations. For instance, “What is 3+2?” might be reformulated as “Calculate the sum of 2 and 3” or “Add two to three”. To avoid train-test leakage, rephrased tasks were consistently placed in the same dataset split. Further details on the annotation process are available in the Appendix.
3.2.5 Reverse Mapping and Filtering
To ensure high-quality data, we incorporate an additional step into the data collection pipeline. We build scripts to map the text-based labels of each bounding box back to their numeric coordinates, and then match the syntax and verify if the task will be executed on the screen. Using this filter, we remove all the non-working or syntactically incorrect data points and finally manually review the set of tasks.
After filtering, we obtain 9802 human-annotated, gold-label data points across more than 200 desktop and web screens (Table 3), split into train, validation, and test sets in a 7:1:2 ratio. All collected data will be publicly released to encourage future work on multimodal agents.
Domain Train Validation Test Total
Mac OS 3028 444 786 4258
Linux 761 126 247 1134
Desktop Windows 1573 216 458 2247
Web - 1427 206 530 2163
Total 6789 992 2,021 9802
Table 3: Dataset distribution across splits and platforms.
Figure 3: DetACT Module. Given an initial image and a natural language task description, we use a pipelined approach to run OCR and SAM on the screen. The outputs from SAM are then used by icon and color-matching modules to obtain an exhaustive set of useful UI elements. The list of elements is passed through LLM based filter to select only the elements related to the given task.
4 Evaluation Metrics
In this section, we detail various evaluation metrics for benchmarking model performance on the OmniACT dataset. UI screens have additional constraints such as spatial relevance which are not factored in most conventional similarity-based metrics such as BLEU [31], CodeBLEU [33], BERTScore [53] and CodeBERTScore [56]. For example, a valid click action is usually not constrained to a single coordinate but can be any coordinate within a specified region. In the event of invalid coordinate predictions, an agent that predicts coordinates further away from the valid region should invoke a higher penalty compared to an agent that predicted coordinates close to the region. We propose two new metrics adapted: Sequence Score (Section 4.1) and Action Score (Section 4.2) aimed at utilizing UI information.
4.1 Sequence Score
The sequence score measures whether the predicted action sequence (e.g., ‘click’, ‘write’, ‘press’) exactly matches the gold sequence. Since predicting the first action in the sequence is relatively straightforward and later actions are more difficult, we define sequence score as follows:
$SeqScore_{i}=\begin{cases}\beta_{1}+\beta_{2}*(s-1)&\text{if all actions match}\\ 0&\text{otherwise}\end{cases}$
where $s$ is the action sequence length, $\beta_{1}$ is set to 0.1 and $\beta_{2}$ is set to 1.
4.2 Action Score
The action score measures how well a code snippet containing the correct action sequence can perform the task. Specifically, for a script with a correct action sequence, we introduce penalties for inaccurate behavior. The penalties are described below:
  1. 1.
    Click penalty ($M$): For actions ‘click’, ‘rightClick’, ‘doubleClick’, ‘moveTo’, and ‘dragTo’, we penalize code snippets where predicted coordinates lie outside of the bounding box of the UI element. The click penalty for the $j^{th}$ action of the $i^{th}$ example is defined as:
[…]
    Here $L_{2}$ corresponds to the smallest Euclidean distance between the predicted coordinate and bounding box. $L_{2}$ is zero when the predicted coordinate lies within the target bounding box. $\mu$ is the Dirichlet smoothing coefficient which we dynamically set to the length of the diagonal of the bounding box. This ensures that the penalty for points outside the bounding box varies based on the size of the bounding box. For two predicted points with the same $L_{2}$, the metric penalizes more heavily if the box is larger. This is sound with the intuition that the chances of clicking on a larger box are higher and should be penalized more in case of a mistake.
  2. 2.
    Key penalty ($K$): For actions ‘press’ and ‘hotkey’, we check whether the set of keys in the target code (represented as $GK_{i}^{j}$) and predicted code (represented as $PK_{i}^{j}$) are the same. It is formally defined as:
[…]
  3. 3.
    Write penalty ($W_{p}$): For action type ‘write’, we penalize the output for the sentence to be typed. Specifically, we the employ BLEU score [31], and compute:
[…]
    Here, $GS_{i}^{j}$ represents the actual sentence to be typed, and $PS_{i}^{j}$ represents the sentence predicted by the model in the $j^{th}$ action of example $i$.
In the above equations, ($\alpha_{i}$) is the weighting factor:
$\alpha_{i}=SeqScore_{i}/\text{ length of sequence $i$}$
This ensures that the action score $\in[0,1]$. The mean action score is calculated as follows:
Action Score = $\textbf{Action Score = }\\ \frac{\sum_{i}max\left(SeqScore_{i}-\sum_{j}(M_{i}^{j}+K_{i}^{j}+W_{i}^{j}),0\right)}{\sum_{i}SeqScore_{i}}$
We report and discuss these metrics for all baseline models in Section 7.
Figure 4: Baseline Model Architecture. Image and task descriptions are sent to DetACT module, which gives a filtered list of UI elements relevant to feed into the prompt along with the task. We also show the prompt structure used for action script generation. This structure is passed through the LLM (along with the image for multimodal LLM) to generate the automation script.
5 DetACT: DETecting ACTions from UI
Understanding UI screens is crucial for multimodal computer tasks. Web-based agents typically use language-only inputs from the HTML DOM. This is insufficient for comprehending the full extent of an application UI, as many components may not be easily described with HTML code. To address this, we propose DetACT, which allows us to convert images of UI layouts into structured code and text outputs for a downstream LLM. DetACT is a system comprised of three distinct modules: the text module, the icon module, and the color module.
  1. 1.
    Text Extraction: We use the EasyOCR model [^2] https://github.com/JaidedAI/EasyOCR to parse over the UI screens and collect all text-based elements. Along with the text, we also note the locations of each of these elements. This is depicted in Figure 3, along with a list of text elements found on the screen using the OCR Module. We segment and classify the different regions within the screenshot using the Segment Anything Model (SAM) [18]. From the outputs, we filter out the non-textual segments for our icon and color detection.
  2. 2.
    Icon Module: For matching with the appropriate icon, we use a pack of 1600 icons [^3] https://icomoon.io/ as templates. Each of these icons is labeled with their appropriate functionality and is matched with the filtered outputs SAM [18]. For the similarity of the two images, we resize the reference icons and segmented region of interest (ROI) to the same size, and convert both images to grayscale. After this, we use the Structural Similarity Index (SSIM) [48], to find the closest match of the ROI to the icons in our set, and select the ones above the SSIM threshold of 0.95. As seen in Figure 3, a few icons matched on the screen are Globe icon, Calendar icon, Person icon, and Location icon; each depicting a different use case.
  3. 3.
    Color Module: Finally, to place all segments of interest into appropriate buckets of colors, we average the RGB pixel values over the ROI and, based on that value, bucket them into different color categories. We categorize colors differently based on the human perspective of the ranges of each color. To avoid ambiguity, we consider eleven major colors, namely yellow, blue, green, red, pink, violet, white, black, orange, brown, and grey. We record the center of the element along with the color.
Once all the elements of each category are extracted with their coordinates, we then filter these UI elements by prompting GPT-4 [30]. We ensure that the elements selected are suited only for our task, for which we also provide the task description in our prompts along with the list of elements. Full details of the prompt are provided in the appendix section of the paper. As we observe in Figure 3, given an image from the Expedia application, and a task (“Click on the Black Location icon and enter the destination as Paris.”), the LLM filters out the elements to retain only “Going To”, “Location Icon”, and the Black colored elements from the screen. This is passed as input to the LLM or vision-language model backbone.
6 Baselines
To evaluate the performance of existing language model-based agents on OmniACT, we conduct experiments with both language-based and multimodal baselines. The DetACT module takes in image and text descriptions of the task and outputs the color, icon, and text-based signals. This is concatenated to the prompt for the LLM prompt-based baselines (see Figure 4). Every prompt starts with a role assignment [55], followed by the detailed API reference of the PyAutoGUI function set, along with a textual description of their function. We then add five in-context examples from the training set that most closely match the task (based on the cosine similarity of the MiniLM [46] embeddings of the reference task and the train examples). We add a list of UI elements filtered by the DetACT module to the prompt. Finally, we provide the rules with the task description. For multimodal baselines, we also pass the image pixels to the vision encoder. We report the results of several baselines:
  • Few-shot Generative LLM: We experiment with models from LLaMA-2 [43], Vicuna-1.5 [7], CodeLLaMA-34B [34], Palmyra [42], and GPT [30] series. We use the prompts structure as shown in Figure 4 to prompt the model. For LLaMA and CodeLLaMa, we reduce the prompt length to 2000 tokens by removing outputs from the DetACT module with lower confidence, as we observed poor performance on longer prompts. For the other models, we allow prompts with up to 4000 token sizes.
  • Finetuned Generative LLM: We fine-tuned the LLaMA-13B model and Vicuna-13B using QLoRa [10] with rank 64 and scaling factor 16 for 300 steps to generate the code given screen description from the DetACT module and the instruction.
  • Few-shot Generative Multimodal Models: As OmniACT is predominantly multimodal, with a majority of tasks being visually grounded, we conduct experiments with large multimodal models. Given the limited research in this domain [52, 47], there is a scarcity of available multimodal models with significant size adept for this task. Here, we experiment with [25, 24], providing a similar prompt as well as the screen image.
We report all results over the test set in Table 4.
7 Results and Analysis
Model
SS($\uparrow$)
$M_{p}$
$K_{p}$
$W_{p}$
AS($\uparrow$)
Prompt based LLMs
LLaMA-7B [43] 4.12 1.24 1.83 0.57 0.48
Vicuna-7B  [7] 3.88 1.17 1.51 0.43 0.77
LLaMA-13B  [43] 4.80 1.32 0.93 0.93 1.62
Vicuna-13B  [7] 5.44 1.65 0.94 1.06 1.78
Palmyra-Instruct-30B [41] 7.51 5.68 0.12 0.40 1.31
CodeLLaMA-34B [35] 10.09 2.99 2.71 0.66 3.72
Palmyra-X 43B [2] 11.20 3.12 3.02 2.12 2.94
GPT-3.5-turbo-0613 [29] 22.85 8.13 4.51 2.31 7.89
GPT-4 [30] 32.75 10.27 6.99 3.89 11.60
Finetuned LLMs
LLaMA-13B FT 8.92 4.61 1.43 0.74 2.14
Vicuna-13B FT 8.78 4.12 1.31 0.63 2.72
Multimodal LLMs
LLaVA-v1.5-7B [25] 13.23 4.73 1.24 1.44 5.82
LLaVA-v1.5-13B [24] 20.56 6.07 3.44 2.85 8.19
Human Performance 82.23 0.12 0.36 1.61 80.14
Table 4: Baseline Performance. (A) Prompt-only LLMs, (B) Fine Tuned LLMs, (C) Prompt-only Multimodal Models. The table represents the Sequence score (SS), click penalty ($M_{p}$), Key penalty ($K_{p}$), Write Penalty ($W_{p}$), and Action Score (AS). The best results for the (SS) and (AS) are highlighted.
Model Sequence Score (normal-$\uparrow$) Action Score (normal-$\uparrow$)
GPT-4 [30] 36.42 12.77
GPT-4V [50] 39.43 20.76
Table 5: Results of GPT-4 and GPT-4V on a subset of 500 samples.
As shown in Table 4, we experiment with three different categories of models, namely Prompt-based LLMs, Fine-tuned LLMs, and Prompt-based Multimodal Models. GPT-4 is the best-performing approach, scoring higher on the sequence score and invoking lower penalties on coordinate predicting and text input. For prompt-only LLMs, the GPT-3.5-turbo and GPT-4 models outperform the other LLM baselines, including the LLaMA [43] and Vicuna [7] models. We observe that CodeLLaMA-34B [35], which is trained for code generation, also achieves a higher performance than other models of the same size at predicting the action sequences.
Fine-tuned models also perform much better than their few-shot prompt-only counterparts. Fine-tuning substantially improves LLaMA-13B’s sequence score (4.80 to 8.92) and action score (1.62 to 2.14), as well as the other metrics. Despite this, we observed that both, prompt-based LLMs and finetuned LLMs face severe mouse penalties, especially on click coordinates. This is because they rely solely on text-based signals.
To address this, we experiment with multimodal language models (Table 4). We observe that the coordinate prediction improves significantly when we provide the entire image as input to the multimodal LLM, as this enables it to fully utilize the screen representation. In addition to open sourced models, we also experiment with the GPT-4-vision-preview API [50] on a subset of 500 data points (due to cost overheads and OpenAI API request limits). Table 5 shows that GPT-4 Vision [50] outperforms GPT-4 significantly on the Action Score along with improving the sequence score, which we attribute to the strong reasoning abilities of GPT-4 coupled with the improved visual understanding capabilities of the GPT-4-vision-preview model [50]. These findings pave the way towards exciting new research directions on building multimodal models for long-horizon planning and code generation.
Human performance over the task: OmniACT consists of visually complicated tasks, and tests various types of computer skills. In order to get a gauge of how well humans perform, we collect evaluation data from human evaluators. We split the test set uniformly amongst 10 human evaluators, and provided them with the screenshot and task instruction. We record the actions taken by the annotators, and measure their performance on our predefined metrics (Table 4). We find that users generally exhibit a high level of proficiency when attempting most tasks for the first time. However, there are instances where users face difficulties in successfully completing certain tasks. These are due to factors including the user’s inability to fully comprehend the task, difficulties in grounding the task to the provided screenshot, or a lack of familiarity with the UI.
8 Conclusion and Future Work
Autonomous virtual agents offer the potential to automate routine tasks, benefiting users with limited technical expertise. To solve this task, we introduce OmniACT, a unique dataset of 9.8K human-labeled data points. OmniACT benchmarks autonomous agents across a range of tasks on web and desktop applications. LLM-based agents, like GPT-4, achieve a respectable action score of 11.6 on our dataset. However, OmniACT presents a challenge for the current state-of-the-art language and multimodal models. It provides a direction for future research on foundational multimodal models that seamlessly integrate language and visual understanding of computer screens and stands poised to drive the next wave of advancements in generalist autonomous agents offering omnipotent assistance to humans.
Appendix
In this paper, we present a novel dataset OmniACT that aids in building more robust multimodal generalist autonomous agents for Desktop and Web. Along with this, we also propose a new continuous scale metric that allows better assessment of actions on a computer screen and the DetACT module which we integrate into multiple LLMs and (Vision Language Models) VLMs that can extract the useful features from the screen image and help us benchmark the dataset. We present the following items that give further insight into the dataset and experiments we performed:
  1. A
    List of applications and websites
  2. B
    Additional Quantitative Results
  3. C
    Annotation Process
  4. D
    Dataset and Metadata Format
  5. E
    Parameters for Model Training
  6. F
    Limitations and Broader Impact
  7. G
    Ethics Statement
  8. H
    Sample of task execution
  9. I
    Prompts for DetACT and Baselines and Sample Responses
Appendix A List of applications and websites
The complete list of applications and websites chosen for the dataset are listed in table 6. We ensure that the data is equitably represented across all domains and operating systems. We choose desktop applications that are more commonly used for each of the operating systems. The commonly used applications we select are representative of eclectic demographics. We also select a few third-party applications that users commonly use on all operating systems. To avoid redundancy and ensure diversity in our tasks, we refrain from choosing the same set of applications for all three operating systems. For websites, we select our candidate set based on user intent. This is sub-categorized into six domains - (1) Shopping, (2) Entertainment, (3) Service, (4) Government, (5) Travel, and (6) Health.
Desktop Websites
MacOS Linux Windows
Mail Music App Store Audible Grammarly AMC Coursera NPS
Maps News Calculator Camera Outlook Apartments FlightAware NY Gov
Message Photos Calendar Desktop Spotify Aramex Google Careers Samsung
Prime Preview Clock Files WeChat Armani HealthLine Trip Advisor
Terminal Reminder Expedia Settings Calendar Asics Hertz UHaul
Weather Stocks Finder Todo Files AT&T Indeed Udemy
Zoom System Preferences iBooks Text Editor Settings BestBuy Instacart United Airlines
Desktop VLC Windows Store Booking Mayo Clinic UPS
Cmu.edu Mint Mobile Walmart
Table 6: List of applications for desktop and web in OmniACT
Appendix B Additional Discussions
B.1 Screen Region Attention
We perform an in-depth analysis of our dataset to see where the coordinate predictions are gravitated. We find interesting insights about the parts of the screen that are attended towards. Most of the predictions for the desktop data are focused either on the center of the screen or the bottom corners. While most of the web interactions are pointing toward the bottom half of the screen, especially concentrating toward the corners. This is depicted in figure 5. This presents some exciting directions for training multimodal models that will heavily rely on the screen image in addition to the elements extracted from DetACT module.
From the histograms at the top and right of the screens in figure 5, we notice that the dataset is not heavily biased towards any particular region of the screen. This implies that models that are likely to attend to only certain regions of the screen or predict the coordinates in the same region of the screen will not have a good performance over the dataset.
Figure 5: Screen region attention in output scripts for desktop and web data
Model
Sequence
Score
Action
Score
GPT-4 (w/o DetACT) 31.87 0.32
GPT-4 + DetACT 32.75 11.60
Table 7: Performance of DetACT Module with GPT-4 Series of Models
B.2 Impact of DetACT Module
In this set of experiments we present the results without using our DetACT module and their comparison with the corresponding model with DetACT outputs. We can clearly observe in table 7 that even models like GPT-4 perform close to zero when not prompted by DetACT as they are only language-based. This highlights the importance of well-equipped screen parsing modules such as our DetACT framework that effectively extract multimodal information in the form of text, which is an essential cue for every model.
B.3 Sequence length versus Performance
We analyze the performance of predicting the sequence of actions as a function of sequence length. Specifically, we want to check whether it becomes difficult for models to predict the sequence correctly when the target sequence is longer. Figure 6 shows the accuracy of sequence prediction as a function of sequence length for GPT-4 on the test set. Only when the entire sequence of actions predicted by the model is the same as the gold sequence of actions, do we give the score as 1. We report the accuracy percentage in the table. From table 6, we observe that the model performance is better for shorter sequences and gets worse as the sequence length increases.
Figure 6: Sequence prediction accuracy of GPT4 on the test set. The performance decreases as the sequence length increases.
Appendix C Annotation Process
C.1 Creating the Bounding Boxes
Figure 7: Annotation Process for creating bounding boxes for desktop applications and websites.
To create the bounding boxes, we use two different techniques for desktop and web data. For desktop data, we use a custom PyQt5 tool that helps us create the bounding boxes using a click-and-drag mechanism. This enhances the bounding box creation process, making it more intuitive and user-friendly. It also enables us to encompass interactive regions of UI elements. For instance, when clicking on a search bar, users can input text, while clicking on a search bar icon (such as a magnifying glass) does not activate the typing area. Next, for the website, we use the DOM from HTML and draw bounding boxes using JavaScript. Both these processes ensure a high degree of accuracy and are done by the authors of the paper. Samples for each of them are shown in figure 7. After the bounding boxes are established, the authors exchange their work and review the boxes created by each other.
C.2 Annotator Guidelines
Figure 8: Screenshots from MTurk Portal that depict the labeling process for all boxes.
For labeling the bounding boxes for both desktop applications and websites, we leverage MTurk. This is depicted in figure 8. MTurk workers are present with a single bounding box on the screen removing the other boxes and are then asked to label the box based on the functionality of the UI element in context to the screen. They are asked to come up with less than a 5-word description or label for each box. Every box is doubly annotated keeping the more appropriate label during the final filtering. Every MTurk worker is paid at an hourly rate of \$20, which is well above Pennsylvania’s minimum wage rate.
Next for formulating the tasks, we recruit student volunteers who have basic knowledge of Python and coding. We train every student volunteer about the PyAutoGUI usage in a thirty-minute session before starting the task formulation process. Every student is then handed over a sample of tasks that can be created using that screen and is asked to come up with more tasks along with their scripts, which are only a few lines. Every volunteer is asked to come up with as many tasks as possible in one hour along with its linguistic variations to ensure diversity in the language of the tasks. Every volunteer is paid an hourly wage rate of \$25, which again surpasses the state’s minimum rate.
We build scripts to take in the tasks generated by the students and run over the screens to ensure the tasks are executable. Once done, the authors perform a final pass to keep only the high-quality tasks.
C.3 Human Feedback and Evaluations
To evaluate the human baselines, we again recruit student volunteers. This time the volunteers are not given the reference output script and only a pair of an image and the task description. The students are asked to then come up with an automation script using a few samples as references. The volunteers use their experience and intuition to comprehend the screen and the functionalities and write scripts based on that. Here the student worker is compensated at an hourly rate of \$25, well exceeding the state’s minimum wage.
Model Hyperparameters
DetACT Filtering GPT-4 temperature = 0.1
LLM Prompt Only Baselines LLaMA-7B max_new_tokens = 400, temperature = 0.8, repetition_penalty = 1.1
Vicuna-7B
LLaMA-13B max_new_tokens = 400, temperature = 0.8
Vicuna-13B
Palmyra-Instruct 30 B temperature = 0.7
CodeLLaMA-34B
max_new_tokens = 400,
temperature = 0.7
Palmyra-X 43B temperature = 0.7
GPT-3.5-turbo temperature = 0.3
GPT-4
LLM Fine-tuned baselines LLaMA-13B FT LoRA rank = 64, scaling factor = 16, LoRA dropout = 0.05, batch size = 4, learning rate = 5e-5
Vicuna-13B FT
Multimodal Model Prompt Only Baselines LLaVA-v1.5-7B temperature = 0.2, max_new_tokens = 512
LLaVA-v1.5-13B
Table 8: Hyperparameters used for each of the models for DetACT filtering and Baselines.
Appendix D Dataset and Metadata Format
The dataset consists of three files, one for each split. Every file has a list of task numbers for that particular split. Every task directory consists of (1) image.png file, which is a screenshot of the screen, and (2) task.txt file which contains the task description and the ground truth output script. We also include another file, (3) box.json, which is part of the metadata and is meant to be only used for running the evaluations that require the coordinates of the boxes and not for performing the task or aiding the model in any way.
Appendix E Parameters for Model Training
We enlist the model specifications for models we run for DetACT filtering and baselines using DetACT in table 8.
Appendix F Broader Impact
The broader impact of our intricately curated dataset is multifaceted. Firstly, our dataset holds the potential to empower the technologically illiterate, providing them with a more automated and accessible way to navigate and operate computers.
Additionally, our dataset opens avenues for groundbreaking UI grounding research, enabling a deeper understanding of user interactions at the operating system level. This, in turn, contributes to advancing the field of artificial intelligence (AI) by expanding the scope of AI to encompass OS-level interactions. The dataset’s rich content further facilitates the enhancement of large language models (LLMs), broadening their skill sets and improving their capabilities in comprehending and responding to diverse user inputs.
This dataset serves as a foundation for the development of multimodal agents capable of operating seamlessly across various screens and executing tasks with long-term horizons. This versatility in application positions the dataset as a valuable resource for the creation of innovative AI-driven solutions that transcend traditional limitations.
An additional and impactful use case involves building assistive tools for handicapped individuals who may face challenges in operating computer systems or performing routine tasks that involve human-computer interaction. By tailoring technologies to meet the needs of individuals with varying abilities, our dataset contributes to the development of assistive technologies that enhance accessibility and improve the quality of life for a diverse range of users.
In summary, the broader impact of our meticulously curated dataset manifests in its potential to drive advancements in technology accessibility, UI grounding research, AI capabilities, and assistive tools, ultimately fostering a more inclusive and empowered technological landscape.
Appendix G Ethics Statement
This work introduces a valuable dataset, yet we recognize a few limitations that exist. State-of-the-art models like GPT-4, may exhibit susceptibility to hallucinations and bias towards specific data types, hindering broad applicability. Reliance on closed models like GPT-4V poses integration challenges due to high costs and time constraints in real-world systems. Despite efforts for equal representation and data collection without personal information, biases may be introduced as the dataset is exclusively in English, and the human-curated content may have temporal biases.
Appendix H Sample of task execution
We present figures 9 containing screenshots and corresponding output scripts for a selected task. In figure 9, the task is taken from a MacOS native application, Books. Here we have a book chapter opened along with the task description to highlight the last line in green color and scroll to the next page. Followed by the task are descriptive screenshots representing every action in the sequence along with its output.
Figure 9: Task Demo along with intermediate steps and corresponding scripts for every action sequence.
Appendix I Prompts for DetACT and Baselines and Sample Responses
In the following pages, we present the full length prompts that we use for the following tasks:
  • Listing 1: Filtering UI elements in DetACT Module
  • Listing 2: Running baselines models using DetACT
Listing 1: Prompt for Filtering UI elements in DetACT Module
Given below are the UI elements extracted from the screen. You have to filter the elements for this UI screen that are relevant for carrying out the task given below.
Make sure to filter the UI elements that may be helpful to carry out the task and mention the element description and the corresponding coordinates for the task.
Format for every element would be in the form of (Element Label, X-Coordinate, Y-Coordinate).
[IMPORTANT!!] Only remove the elements if you are sure that they will not help the task.
[IMPORTANT!!] Follow the output structure strictly as given in the below example and output nothing else other than the required output.
Sample Task:
{SAMPLE_TASK}
Sample UI Elements:
{LIST_OF_SAMPLE_UI_ELEMENTS}
Sample Filtered UI Elements:
{LIST_OF_SAMPLE_FILTERED_UI_ELEMENTS}
Given Task:
{TASK}
Given UI Elements:
{LIST_OF_UI_ELEMENTS}
Listing 2: Prompt for baselines models using DetACT outputs
You are an excellent robotic process automation agent who needs to generate a pyautogui script for the tasks given to you. You will be given UI elements on the screen that you can interact with as a list of tuples, where each tuple will be of the form [UI Text, X coordinate, Y coordinate]. You will also be given an example to help with the format of the script that needs to be generated.
[IMPORTANT!!] Stick to the format of the Output scripts in the example
[IMPORTANT!!] Use only the functions from API docs
[IMPORTANT!!] Follow the Output format strictly. Only write the script and nothing else.
Here is the API reference for generating script:
def click(x=moveToX, y=moveToY):
    """takes the mouse to location (moveToX, moveToY) and does a left click
    Example:
    High Level Goal: Click at co-ordinate (150, 230).
    Python script:
    import pyautogui
    pyautogui.click(150, 230)
    pass
def rightClick(x=moveToX, y=moveToY):
    """takes the mouse to location (moveToX, moveToY) and does a right click
    Example:
    High Level Goal: Right click at co-ordiante (350, 680).
    Python script:
    import pyautogui
    pyautogui.rightClick(350, 680)
    pass
def doubleClick(x=moveToX, y=moveToY):
    """takes the mouse to location (moveToX, moveToY) and does a right click
    Example:
    High Level Goal: Right click at co-ordiante (350, 680).
    Python script:
    import pyautogui
    pyautogui.rightClick(350, 680)
    pass
def scroll(clicks=amount_to_scroll):
    """scrolls the window that has mouse pointer by float value (amount_to_scroll)
    Example:
    High Level Goal: Scroll screen by (30).
    Python script:
    import pyautogui
    pyautogui.scroll(30)
    pass
def hscroll(clicks=amount_to_scroll):
    """scrolls the window that has mouse pointer horizontally by float value (amount_to_scroll)
    Example:
    High Level Goal: Scroll screen horizontally by (30).
    Python script:
    import pyautogui
    pyautogui.hscroll(30)
    pass
def dragTo(x=moveToX, y=moveToY, button=holdButton):
    """drags the mouse to (moveToX, moveToY) with (holdButton) pressed. holdButton can be ’left’, ’middle’ or ’right’.
    Example:
    High Level Goal: drag the screen from current position to (450, 600) with left click of the mouse.
    Python script:
    import pyautogui
    pyautogui.dragTo(450, 600, ’left’)
    pass
def pyautogui.moveTo(x=moveToX, y=moveToY)
    """take the mouse pointer to (moveToX, moveToY)
    Example:
    High Level Goal: hover the mouse pointer to (450, 600).
    Python script:
    import pyautogui
    pyautogui.moveTo(450, 600)
    pass
def write(str=stringType, interval=secs_between_keys):
    """writes the string wherever keyboard cursor is at the function calling time with (secs_between_keys) seconds between characters
    Example:
    High Level Goal: Write "Hello world" with 0.1 seconds rate
    Python script:
    import pyautogui
    pyautogui.write("Hello world", 0.1)
    pass
def press(str=string_to_type):
    """simulate pressing a key down and then releasing it up. Sample keys include ’enter’, ’shift’, arrow keys, ’f1’.
    Example:
    High Level Goal: Press enter key now
    Python script:
    import pyautogui
    pyautogui.press("enter")
    pass
def hotkey(*args = list_of_hotkey):
    """Keyboard hotkeys like Ctrl-S or Ctrl-Shift-1 can be done by passing a list of key names to hotkey(). Multiple keys can be pressed together with a hotkey.
    Example:
    High Level Goal: Use Ctrl and V to paste from clipboard
    Python script:
    import pyautogui
    pyautogui.hotkey("ctrl", "v")
    pass
Here are a few examples for generating the script and format to be followed:
Example 1:
{RETREIVED_EXAMPLE_1}
Example 2:
{RETREIVED_EXAMPLE_2}
Example 3:
{RETREIVED_EXAMPLE_3}
Example 4:
{RETREIVED_EXAMPLE_4}
Example 5:
{RETREIVED_EXAMPLE_5}
Here are the Textual Elements found on the screen along with their coordinates:
{TEXTUAL_ELEMENTS_FROM_DETACT}
Here are the Icon/Image Elements found on the screen along with their coordinates:
{ICON_ELEMENTS_FROM_DETACT}
Here are the Color Elements found on the screen along with their coordinates:
{COLOR_ELEMENTS_FROM_DETACT}
Based on the list of UI elements and the examples above, generate the pyautogui script for the following task:
{TASK}

### v3

@@
-OmniACT: A Dataset and Benchmark for Enabling Multimodal Generalist Autonomous Agents for Desktop and Web
+https://huggingface.co/datasets/Writer/omniact
 Abstract
-footnotetext: $\dagger$ These authors contributed equally. The order is determined by dice rolling.
+footnotetext: * These authors contributed equally. The order is determined by dice rolling.
 For decades, human-computer interaction has fundamentally been manual.
@@
 Our scope extends beyond traditional web automation, covering a diverse range of desktop applications.
-The dataset consists of fundamental tasks such as “Play the next song”, as well as longer horizon tasks such as “Send an email to John Doe mentioning the time and place to meet”.
+The dataset consists of fundamental tasks such as “Play the next song", as well as longer horizon tasks such as “Send an email to John Doe mentioning the time and place to meet".
 Specifically, given a pair of screen image and a visually-grounded natural language task, the goal is to generate a script capable of fully executing the task.
@@
 Our benchmark provides a platform to measure and evaluate the progress of language model agents in automating computer tasks and motivates future work towards building multimodal models that bridge large language models and the visual grounding of computer screens.
+Keywords:
+Generalist Agent Multimodal Machine Learning Dataset and Benchmark Vision Language Understanding UI grounding Human-computer interaction
 Figure 1: OmniACT dataset and benchmark for enabling autonomous human-computer interaction agents.
 The left shows an image paired with a natural language task description as the input, the right shows the resulting action script to be executed on the screen.
-Examples are presented from Stocks, Apartments.com, and Weather application.
+Examples are presented from Stocks, Apartments, and Weather application.
 1 Introduction
 Performing computer tasks based on natural language instructions has been a long-standing goal of artificial intelligence [49].
-One concrete objective in the line of research is to develop generalist agents that can assist humans in doing computer tasks [19], such as “Order a pizza from Domino’s” or “Write a message to John.” The agent should be able to open the application and perform the task.
+One concrete objective in the line of research is to develop generalist agents that can assist humans in doing computer tasks [21], such as “Order a pizza from Domino’s" or “Write a message to John." The agent should be able to open the application and perform the task.
 Executing these actions on a personal computer involves a sequence of interactions with a mouse and keyboard.
@@
 We experiment with fine-tuning Vicuna-13B and LLaMA-13B models using QLoRA [10].
-We also benchmark multimodal baseline LLaVa-v1.5-7B, LLaVa-v1.5-13B [43] and GPT-4-vision-preview [50] for the task.
+We also benchmark multimodal baseline LLaVa-v1.5-7B, LLaVa-v1.5-13B [47], Gemini-Pro [44] and GPT-4-vision-preview [55] for the task.
 Our findings highlight the necessity for a multimodal model capable of executing these tasks, and our analysis provides insights into promising future work in the space.
@@
 2 Related Work
+Table 1: Comparison of OmniACT with other related benchmarks.
 Datasets Size Env Type
@@
 Task
+VisualWebArena [20] 910 Web Yes Yes Yes No No Web Navigation
 WebArena [62] 812 Web Yes Yes Yes No No Web Navigation
@@
 OmniACT (Ours) 9802 Desktop and Web Yes Yes Yes Yes Yes Code Generation
-Table 1: Comparison of OmniACT with other related benchmarks.
 2.1 UI Understanding
@@
 The majority of machine learning models trained for UI understanding leverage the Rico dataset [8] and its extensions, which contain 64,462 unique Android screens and metadata.
-In addition, Banerjee et al.
-[4] released the UICaptions dataset, which consists of diverse image-captions pairs across a wide range of applications.
-PixelHelp [21] also released a corpus to train models that can interpret natural language instructions and map them to mobile user interface actions.
+In addition, [4] released the UICaptions dataset, which consists of diverse image-captions pairs across a wide range of applications.
+PixelHelp [23] also released a corpus to train models that can interpret natural language instructions and map them to mobile UI actions.
 2.2 Autonomous Computer Agents
@@
 Recent research such as ViperGPT [43] Chameleon [29], RCI Agent [18], VisProg [12], and [31] employ LLMs for planning or action prediction in developing autonomous agents.
-Benchmark datasets, such as MiniWoB [37], WebShop [51], Macaw-LLM [27], ASH-Prompting [38] Mind2Web [9] and WebArena [57] have also been proposed to measure the ability of LLM-based agents in automating web tasks.
+Benchmark datasets, such as MiniWoB [40], WebShop [56], Macaw-LLM [30], ASH-Prompting [41] Mind2Web [9], WebArena [62], AgentBench [28] and VisualWebArena [20] have also been proposed to measure the ability of LLM-based agents to automate web tasks.
 These methods mainly involve agents that operate on a text-based Document Object Model (DOM) of HTML scripts.
 This limits their understanding of screen context, which is crucial for the model’s decision-making and action-taking processes.
-To address this limitation, Rawles et al.
-[32] released Android in the Wild, a dataset comprising screens, natural language instructions, and corresponding actions.
+To address this limitation, [35] released Android in the Wild, a dataset comprising screens, natural language instructions, and corresponding actions.
 Following this,  [59] proposed a multimodal model, AutoUI, which is designed to build an agent on the Android in the Wild dataset confined to the Android ecosystem.
+WebAgent [13] utilized Flan-U-PaLM, for grounded code generation, and HTML-T5 and showed improvement on real-world websites.
 Current benchmarks for autonomous agents focus mainly on the Web or Android environments, posing challenges for tasks involving desktop applications or spanning multiple applications beyond the web domain.
@@
 To execute these tasks, a user must accurately perform a series of operations using the mouse and keyboard.
-For example, to check the rate of change in Google’s stock price over the last month, the mouse has to be moved to the last month and dragged while holding the left-click button to the current month.
+Eg: to check the rate of change in Google’s stock price over the last month, the mouse has to be moved to the last month and dragged while holding the left-click button to the current month.
 3.2 Dataset Preparation
@@
 This ensured the high quality and functionality of our dataset, making it a valuable resource for training and evaluating autonomous agents.
+Table 2: Action types supported by OmniACT and the number of instances for each action in the dataset.
 Type Action %
@@
 Keyboard Write 11.65
-Table 2: Action type supported by OmniACT and the number of instances for each of the actions in the dataset.
 3.2.1 Application/Website Selection
@@
     Desktop Applications: We build a custom annotation interface based on PyQt5111https://pypi.org/project/PyQt5/ to create bounding boxes manually over a screen image using a simple drag-and-click mechanism.
-This custom interface expedites the process and allows us to get highly accurate gold-label data points for desktop images.
+This custom interface expedites the process and allows us to get highly accurate gold-label data for desktop images.
   2. 2.
-    Websites: For webpages, we write JavaScript code to extract all interactable (click, hover, type, etc.) regions from the HTML source code.
+    Websites: For webpages, we write JavaScript code to extract all interactable (click, type, etc.) regions from HTML source code.
 We also extract banners, dropdowns, submit, and radio buttons from the screen.
@@
 To map each bounding box to its correct functional description, we leverage Amazon MTurk workers (see details in Appendix), who are given an image with a bounding box and are required to write the correct description or label of the bounding box’s function.
-For example, given an image of an Amazon webpage with a search bar, the annotator labels it as “find-product-search-bar”.
+For example, given an image of an Amazon webpage with a search bar, the annotator labels it as “find-product-search-bar".
 The logical descriptions are used to create tasks in a structured manner without the need to identify individual bounding box coordinates.
 3.2.4 Task Creation
+Figure 3: DetACT Module.
+Given an initial image and a natural language task description, we use a pipelined approach to run OCR and SAM on the screen.
+The outputs from SAM are then used by icon and color-matching modules to obtain an exhaustive set of useful UI elements.
+The list of elements is passed through LLM based filter to select only the elements related to the given task.
 Our approach for each screen involves utilizing all human-annotated bounding boxes and their labels to create tasks that can be executed within the confines of a single screen.
@@
 Each student generated multiple tasks, each accompanied by three alternative natural language reformulations.
-For instance, “What is 3+2?” might be reformulated as “Calculate the sum of 2 and 3” or “Add two to three”.
+For instance, “What is 3+2?" might be reformulated as “Calculate the sum of 2 and 3" or “Add two to three".
 To avoid train-test leakage, rephrased tasks were consistently placed in the same dataset split.
@@
 3.2.5 Reverse Mapping and Filtering
-To ensure high-quality data, we incorporate an additional step into the data collection pipeline.
-We build scripts to map the text-based labels of each bounding box back to their numeric coordinates, and then match the syntax and verify if the task will be executed on the screen.
-Using this filter, we remove all the non-working or syntactically incorrect data points and finally manually review the set of tasks.
-After filtering, we obtain 9802 human-annotated, gold-label data points across more than 200 desktop and web screens (Table 3), split into train, validation, and test sets in a 7:1:2 ratio.
-All collected data will be publicly released to encourage future work on multimodal agents.
+Table 3: Dataset distribution across splits and platforms.
 Domain Train Validation Test Total
@@
 Total 6789 992 2,021 9802
-Table 3: Dataset distribution across splits and platforms.
-Figure 3: DetACT Module.
-Given an initial image and a natural language task description, we use a pipelined approach to run OCR and SAM on the screen.
-The outputs from SAM are then used by icon and color-matching modules to obtain an exhaustive set of useful UI elements.
-The list of elements is passed through LLM based filter to select only the elements related to the given task.
+To ensure high-quality data, we incorporate an additional step into the data collection pipeline.
+We build scripts to map the text-based labels of each bounding box back to their numeric coordinates, and then match the syntax and verify if the task will be executed on the screen.
+Using this filter, we remove all the non-working or syntactically incorrect data points and finally manually review the set of tasks.
+After filtering, we obtain 9802 human-annotated, gold-label data points across more than 200 desktop and web screens (Table 3), split into train, validation, and test sets in a 7:1:2 ratio.
+All collected data will be publicly released to encourage future work on multimodal agents.
 4 Evaluation Metrics
@@
 Since predicting the first action in the sequence is relatively straightforward and later actions are more difficult, we define sequence score as follows:
-$SeqScore_{i}=\begin{cases}\beta_{1}+\beta_{2}*(s-1)&\text{if all actions match}\\ 0&\text{otherwise}\end{cases}$
+$SeqScore_{i}=\begin{cases}\beta_{1}+\beta_{2}*(s-1)&\text{if all actions match}\\ 0&\text{otherwise}\end{cases}$ (1)
 where $s$ is the action sequence length, $\beta_{1}$ is set to 0.1 and $\beta_{2}$ is set to 1.
@@
 $L_{2}$ is zero when the predicted coordinate lies within the target bounding box.
-$\mu$ is the Dirichlet smoothing coefficient which we dynamically set to the length of the diagonal of the bounding box.
+$\mu$ is the Dirichlet smoothing coefficient which we dynamically set to the inverse of the length of the diagonal of the bounding box.
 This ensures that the penalty for points outside the bounding box varies based on the size of the bounding box.
@@
 In the above equations, ($\alpha_{i}$) is the weighting factor:
-$\alpha_{i}=SeqScore_{i}/\text{ length of sequence $i$}$
+$\alpha_{i}=SeqScore_{i}/\text{ length of sequence $i$}$ (5)
 This ensures that the action score $\in[0,1]$. The mean action score is calculated as follows:
-Action Score = $\textbf{Action Score = }\\ \frac{\sum_{i}max\left(SeqScore_{i}-\sum_{j}(M_{i}^{j}+K_{i}^{j}+W_{i}^{j}),0\right)}{\sum_{i}SeqScore_{i}}$
-We report and discuss these metrics for all baseline models in Section 7.
+Action Score = $\textbf{Action Score = }\frac{\sum_{i}max\left(SeqScore_{i}-\sum_{j}(M_{i}^{j}+K_{i}^{j}+W_{i}^{j}),0\right)}{\sum_{i}SeqScore_{i}}$ (6)
+5 DetACT: DETecting ACTions from UI
 Figure 4: Baseline Model Architecture.
@@
 This structure is passed through the LLM (along with the image for multimodal LLM) to generate the automation script.
-5 DetACT: DETecting ACTions from UI
 Understanding UI screens is crucial for multimodal computer tasks.
@@
 Full details of the prompt are provided in the appendix section of the paper.
-As we observe in Figure 3, given an image from the Expedia application, and a task (“Click on the Black Location icon and enter the destination as Paris.”), the LLM filters out the elements to retain only “Going To”, “Location Icon”, and the Black colored elements from the screen.
-This is passed as input to the LLM or vision-language model backbone.
+As we observe in Figure 3, given an image from the Expedia application, and a task (“Click on the Black Location icon and enter the destination as Paris."), the LLM filters out the elements to retain only “Going To", “Location Icon", and the Black colored elements from the screen.
+This is passed as input to the LLM or VLM backbone.
 6 Baselines
-To evaluate the performance of existing language model-based agents on OmniACT, we conduct experiments with both language-based and multimodal baselines.
+To evaluate the performance of existing language model-based agents, we conduct experiments with both language-based and multimodal baselines.
 The DetACT module takes in image and text descriptions of the task and outputs the color, icon, and text-based signals.
@@
 For multimodal baselines, we also pass the image pixels to the vision encoder.
+We choose coordinate-based UI elements in the prompt as recent techniques like the Set-of-Mark (SOM) [54] prompting does not work for desktop settings since it is difficult to obtain interactive elements from the desktop screen images.
 We report the results of several baselines:
@@
 Here, we experiment with [27, 26], providing a similar prompt as well as the screen image.
-We report all results over the test set in Table 4.
 7 Results and Analysis
+Table 4: Baseline Performance.
+(A) Prompt-only LLMs, (B) Fine Tuned LLMs, (C) Prompt-only Multimodal Models.
+The table represents the Sequence score (SS), click penalty ($M_{p}$), Key penalty ($K_{p}$), Write Penalty ($W_{p}$), and Action Score (AS).
+The best results for the (SS) and (AS) are highlighted.
 Model
@@
 LLaVA-v1.5-13B [26] 20.56 6.07 3.44 2.85 8.19
+Gemini-Pro [44] 30.98 9.05 6.81 3.66 11.46
+GPT-4V [26] 38.72 10.53 7.14 4.03 17.02
 Human Performance 82.23 0.12 0.36 1.61 80.14
-Table 4: Baseline Performance.
-(A) Prompt-only LLMs, (B) Fine Tuned LLMs, (C) Prompt-only Multimodal Models.
-The table represents the Sequence score (SS), click penalty ($M_{p}$), Key penalty ($K_{p}$), Write Penalty ($W_{p}$), and Action Score (AS).
-The best results for the (SS) and (AS) are highlighted.
-Model Sequence Score (normal-$\uparrow$) Action Score (normal-$\uparrow$)
-GPT-4 [30] 36.42 12.77
-GPT-4V [50] 39.43 20.76
-Table 5: Results of GPT-4 and GPT-4V on a subset of 500 samples.
 As shown in Table 4, we experiment with three different categories of models, namely Prompt-based LLMs, Fine-tuned LLMs, and Prompt-based Multimodal Models.
@@
 We observe that the coordinate prediction improves significantly when we provide the entire image as input to the multimodal LLM, as this enables it to fully utilize the screen representation.
-In addition to open sourced models, we also experiment with the GPT-4-vision-preview API [50] on a subset of 500 data points (due to cost overheads and OpenAI API request limits).
-Table 5 shows that GPT-4 Vision [50] outperforms GPT-4 significantly on the Action Score along with improving the sequence score, which we attribute to the strong reasoning abilities of GPT-4 coupled with the improved visual understanding capabilities of the GPT-4-vision-preview model [50].
+In addition to open sourced models, we also experiment with the GPT-4-vision API [55] which shows that GPT-4 Vision [55] outperforms GPT-4 significantly on the Action Score along with improving the sequence score, which we attribute to the strong reasoning abilities of GPT-4 coupled with the improved visual understanding capabilities of the GPT-4-vision model [55].
 These findings pave the way towards exciting new research directions on building multimodal models for long-horizon planning and code generation.
@@
 It provides a direction for future research on foundational multimodal models that seamlessly integrate language and visual understanding of computer screens and stands poised to drive the next wave of advancements in generalist autonomous agents offering omnipotent assistance to humans.
-Appendix
-In this paper, we present a novel dataset OmniACT that aids in building more robust multimodal generalist autonomous agents for Desktop and Web.
-Along with this, we also propose a new continuous scale metric that allows better assessment of actions on a computer screen and the DetACT module which we integrate into multiple LLMs and (Vision Language Models) VLMs that can extract the useful features from the screen image and help us benchmark the dataset.
-We present the following items that give further insight into the dataset and experiments we performed:
-  1. A
-    List of applications and websites
-  2. B
-    Additional Quantitative Results
-  3. C
-    Annotation Process
-  4. D
-    Dataset and Metadata Format
-  5. E
-    Parameters for Model Training
-  6. F
-    Limitations and Broader Impact
-  7. G
-    Ethics Statement
-  8. H
-    Sample of task execution
-  9. I
-    Prompts for DetACT and Baselines and Sample Responses
-Appendix A List of applications and websites
-The complete list of applications and websites chosen for the dataset are listed in table 6.
-We ensure that the data is equitably represented across all domains and operating systems.
-We choose desktop applications that are more commonly used for each of the operating systems.
-The commonly used applications we select are representative of eclectic demographics.
-We also select a few third-party applications that users commonly use on all operating systems.
-To avoid redundancy and ensure diversity in our tasks, we refrain from choosing the same set of applications for all three operating systems.
-For websites, we select our candidate set based on user intent.
-This is sub-categorized into six domains - (1) Shopping, (2) Entertainment, (3) Service, (4) Government, (5) Travel, and (6) Health.
-Desktop Websites
-MacOS Linux Windows
-Mail Music App Store Audible Grammarly AMC Coursera NPS
-Maps News Calculator Camera Outlook Apartments FlightAware NY Gov
-Message Photos Calendar Desktop Spotify Aramex Google Careers Samsung
-Prime Preview Clock Files WeChat Armani HealthLine Trip Advisor
-Terminal Reminder Expedia Settings Calendar Asics Hertz UHaul
-Weather Stocks Finder Todo Files AT&T Indeed Udemy
-Zoom System Preferences iBooks Text Editor Settings BestBuy Instacart United Airlines
-Desktop VLC Windows Store Booking Mayo Clinic UPS
-Cmu.edu Mint Mobile Walmart
-Table 6: List of applications for desktop and web in OmniACT
-Appendix B Additional Discussions
-B.1 Screen Region Attention
-We perform an in-depth analysis of our dataset to see where the coordinate predictions are gravitated.
-We find interesting insights about the parts of the screen that are attended towards.
-Most of the predictions for the desktop data are focused either on the center of the screen or the bottom corners.
-While most of the web interactions are pointing toward the bottom half of the screen, especially concentrating toward the corners.
-This is depicted in figure 5.
-This presents some exciting directions for training multimodal models that will heavily rely on the screen image in addition to the elements extracted from DetACT module.
-From the histograms at the top and right of the screens in figure 5, we notice that the dataset is not heavily biased towards any particular region of the screen.
-This implies that models that are likely to attend to only certain regions of the screen or predict the coordinates in the same region of the screen will not have a good performance over the dataset.
-Figure 5: Screen region attention in output scripts for desktop and web data
-Model
-Sequence
-Score
-Action
-Score
-GPT-4 (w/o DetACT) 31.87 0.32
-GPT-4 + DetACT 32.75 11.60
-Table 7: Performance of DetACT Module with GPT-4 Series of Models
-B.2 Impact of DetACT Module
-In this set of experiments we present the results without using our DetACT module and their comparison with the corresponding model with DetACT outputs.
-We can clearly observe in table 7 that even models like GPT-4 perform close to zero when not prompted by DetACT as they are only language-based.
-This highlights the importance of well-equipped screen parsing modules such as our DetACT framework that effectively extract multimodal information in the form of text, which is an essential cue for every model.
-B.3 Sequence length versus Performance
-We analyze the performance of predicting the sequence of actions as a function of sequence length.
-Specifically, we want to check whether it becomes difficult for models to predict the sequence correctly when the target sequence is longer.
-Figure 6 shows the accuracy of sequence prediction as a function of sequence length for GPT-4 on the test set.
-Only when the entire sequence of actions predicted by the model is the same as the gold sequence of actions, do we give the score as 1.
-We report the accuracy percentage in the table.
-From table 6, we observe that the model performance is better for shorter sequences and gets worse as the sequence length increases.
-Figure 6: Sequence prediction accuracy of GPT4 on the test set. The performance decreases as the sequence length increases.
-Appendix C Annotation Process
-C.1 Creating the Bounding Boxes
-Figure 7: Annotation Process for creating bounding boxes for desktop applications and websites.
-To create the bounding boxes, we use two different techniques for desktop and web data.
-For desktop data, we use a custom PyQt5 tool that helps us create the bounding boxes using a click-and-drag mechanism.
-This enhances the bounding box creation process, making it more intuitive and user-friendly.
-It also enables us to encompass interactive regions of UI elements.
-For instance, when clicking on a search bar, users can input text, while clicking on a search bar icon (such as a magnifying glass) does not activate the typing area.
-Next, for the website, we use the DOM from HTML and draw bounding boxes using JavaScript.
-Both these processes ensure a high degree of accuracy and are done by the authors of the paper.
-Samples for each of them are shown in figure 7.
-After the bounding boxes are established, the authors exchange their work and review the boxes created by each other.
-C.2 Annotator Guidelines
-Figure 8: Screenshots from MTurk Portal that depict the labeling process for all boxes.
-For labeling the bounding boxes for both desktop applications and websites, we leverage MTurk.
-This is depicted in figure 8.
-MTurk workers are present with a single bounding box on the screen removing the other boxes and are then asked to label the box based on the functionality of the UI element in context to the screen.
-They are asked to come up with less than a 5-word description or label for each box.
-Every box is doubly annotated keeping the more appropriate label during the final filtering.
-Every MTurk worker is paid at an hourly rate of \$20, which is well above Pennsylvania’s minimum wage rate.
-Next for formulating the tasks, we recruit student volunteers who have basic knowledge of Python and coding.
-We train every student volunteer about the PyAutoGUI usage in a thirty-minute session before starting the task formulation process.
-Every student is then handed over a sample of tasks that can be created using that screen and is asked to come up with more tasks along with their scripts, which are only a few lines.
-Every volunteer is asked to come up with as many tasks as possible in one hour along with its linguistic variations to ensure diversity in the language of the tasks.
-Every volunteer is paid an hourly wage rate of \$25, which again surpasses the state’s minimum rate.
-We build scripts to take in the tasks generated by the students and run over the screens to ensure the tasks are executable.
-Once done, the authors perform a final pass to keep only the high-quality tasks.
-C.3 Human Feedback and Evaluations
-To evaluate the human baselines, we again recruit student volunteers.
-This time the volunteers are not given the reference output script and only a pair of an image and the task description.
-The students are asked to then come up with an automation script using a few samples as references.
-The volunteers use their experience and intuition to comprehend the screen and the functionalities and write scripts based on that.
-Here the student worker is compensated at an hourly rate of \$25, well exceeding the state’s minimum wage.
-Model Hyperparameters
-DetACT Filtering GPT-4 temperature = 0.1
-LLM Prompt Only Baselines LLaMA-7B max_new_tokens = 400, temperature = 0.8, repetition_penalty = 1.1
-Vicuna-7B
-LLaMA-13B max_new_tokens = 400, temperature = 0.8
-Vicuna-13B
-Palmyra-Instruct 30 B temperature = 0.7
-CodeLLaMA-34B
-max_new_tokens = 400,
-temperature = 0.7
-Palmyra-X 43B temperature = 0.7
-GPT-3.5-turbo temperature = 0.3
-GPT-4
-LLM Fine-tuned baselines LLaMA-13B FT LoRA rank = 64, scaling factor = 16, LoRA dropout = 0.05, batch size = 4, learning rate = 5e-5
-Vicuna-13B FT
-Multimodal Model Prompt Only Baselines LLaVA-v1.5-7B temperature = 0.2, max_new_tokens = 512
-LLaVA-v1.5-13B
-Table 8: Hyperparameters used for each of the models for DetACT filtering and Baselines.
-Appendix D Dataset and Metadata Format
-The dataset consists of three files, one for each split.
-Every file has a list of task numbers for that particular split.
-Every task directory consists of (1) image.png file, which is a screenshot of the screen, and (2) task.txt file which contains the task description and the ground truth output script.
-We also include another file, (3) box.json, which is part of the metadata and is meant to be only used for running the evaluations that require the coordinates of the boxes and not for performing the task or aiding the model in any way.
-Appendix E Parameters for Model Training
-We enlist the model specifications for models we run for DetACT filtering and baselines using DetACT in table 8.
-Appendix F Broader Impact
-The broader impact of our intricately curated dataset is multifaceted.
-Firstly, our dataset holds the potential to empower the technologically illiterate, providing them with a more automated and accessible way to navigate and operate computers.
-Additionally, our dataset opens avenues for groundbreaking UI grounding research, enabling a deeper understanding of user interactions at the operating system level.
-This, in turn, contributes to advancing the field of artificial intelligence (AI) by expanding the scope of AI to encompass OS-level interactions.
-The dataset’s rich content further facilitates the enhancement of large language models (LLMs), broadening their skill sets and improving their capabilities in comprehending and responding to diverse user inputs.
-This dataset serves as a foundation for the development of multimodal agents capable of operating seamlessly across various screens and executing tasks with long-term horizons.
-This versatility in application positions the dataset as a valuable resource for the creation of innovative AI-driven solutions that transcend traditional limitations.
-An additional and impactful use case involves building assistive tools for handicapped individuals who may face challenges in operating computer systems or performing routine tasks that involve human-computer interaction.
-By tailoring technologies to meet the needs of individuals with varying abilities, our dataset contributes to the development of assistive technologies that enhance accessibility and improve the quality of life for a diverse range of users.
-In summary, the broader impact of our meticulously curated dataset manifests in its potential to drive advancements in technology accessibility, UI grounding research, AI capabilities, and assistive tools, ultimately fostering a more inclusive and empowered technological landscape.
-Appendix G Ethics Statement
+9 Limitations
 This work introduces a valuable dataset, yet we recognize a few limitations that exist.
 State-of-the-art models like GPT-4, may exhibit susceptibility to hallucinations and bias towards specific data types, hindering broad applicability.
-Reliance on closed models like GPT-4V poses integration challenges due to high costs and time constraints in real-world systems.
-Despite efforts for equal representation and data collection without personal information, biases may be introduced as the dataset is exclusively in English, and the human-curated content may have temporal biases.
-Appendix H Sample of task execution
-We present figures 9 containing screenshots and corresponding output scripts for a selected task.
-In figure 9, the task is taken from a MacOS native application, Books.
-Here we have a book chapter opened along with the task description to highlight the last line in green color and scroll to the next page.
-Followed by the task are descriptive screenshots representing every action in the sequence along with its output.
-Figure 9: Task Demo along with intermediate steps and corresponding scripts for every action sequence.
-Appendix I Prompts for DetACT and Baselines and Sample Responses
-In the following pages, we present the full length prompts that we use for the following tasks:
-  • Listing 1: Filtering UI elements in DetACT Module
-  • Listing 2: Running baselines models using DetACT
-Listing 1: Prompt for Filtering UI elements in DetACT Module
-Given below are the UI elements extracted from the screen. You have to filter the elements for this UI screen that are relevant for carrying out the task given below.
-Make sure to filter the UI elements that may be helpful to carry out the task and mention the element description and the corresponding coordinates for the task.
-Format for every element would be in the form of (Element Label, X-Coordinate, Y-Coordinate).
-[IMPORTANT!!] Only remove the elements if you are sure that they will not help the task.
-[IMPORTANT!!] Follow the output structure strictly as given in the below example and output nothing else other than the required output.
-Sample Task:
-{SAMPLE_TASK}
-Sample UI Elements:
-{LIST_OF_SAMPLE_UI_ELEMENTS}
-Sample Filtered UI Elements:
-{LIST_OF_SAMPLE_FILTERED_UI_ELEMENTS}
-Given Task:
-{TASK}
-Given UI Elements:
-{LIST_OF_UI_ELEMENTS}
-Listing 2: Prompt for baselines models using DetACT outputs
-You are an excellent robotic process automation agent who needs to generate a pyautogui script for the tasks given to you.
-You will be given UI elements on the screen that you can interact with as a list of tuples, where each tuple will be of the form [UI Text, X coordinate, Y coordinate].
-You will also be given an example to help with the format of the script that needs to be generated.
-[IMPORTANT!!] Stick to the format of the Output scripts in the example
-[IMPORTANT!!] Use only the functions from API docs
-[IMPORTANT!!] Follow the Output format strictly. Only write the script and nothing else.
-Here is the API reference for generating script:
-def click(x=moveToX, y=moveToY):
-    """takes the mouse to location (moveToX, moveToY) and does a left click
-    Example:
-    High Level Goal: Click at co-ordinate (150, 230).
-    Python script:
-    import pyautogui
-    pyautogui.click(150, 230)
-    pass
-def rightClick(x=moveToX, y=moveToY):
-    """takes the mouse to location (moveToX, moveToY) and does a right click
-    Example:
-    High Level Goal: Right click at co-ordiante (350, 680).
-    Python script:
-    import pyautogui
-    pyautogui.rightClick(350, 680)
-    pass
-def doubleClick(x=moveToX, y=moveToY):
-    """takes the mouse to location (moveToX, moveToY) and does a right click
-    Example:
-    High Level Goal: Right click at co-ordiante (350, 680).
-    Python script:
-    import pyautogui
-    pyautogui.rightClick(350, 680)
-    pass
-def scroll(clicks=amount_to_scroll):
-    """scrolls the window that has mouse pointer by float value (amount_to_scroll)
-    Example:
-    High Level Goal: Scroll screen by (30).
-    Python script:
-    import pyautogui
-    pyautogui.scroll(30)
-    pass
-def hscroll(clicks=amount_to_scroll):
-    """scrolls the window that has mouse pointer horizontally by float value (amount_to_scroll)
-    Example:
-    High Level Goal: Scroll screen horizontally by (30).
-    Python script:
-    import pyautogui
-    pyautogui.hscroll(30)
-    pass
-def dragTo(x=moveToX, y=moveToY, button=holdButton):
-    """drags the mouse to (moveToX, moveToY) with (holdButton) pressed. holdButton can be ’left’, ’middle’ or ’right’.
-    Example:
-    High Level Goal: drag the screen from current position to (450, 600) with left click of the mouse.
-    Python script:
-    import pyautogui
-    pyautogui.dragTo(450, 600, ’left’)
-    pass
-def pyautogui.moveTo(x=moveToX, y=moveToY)
-    """take the mouse pointer to (moveToX, moveToY)
-    Example:
-    High Level Goal: hover the mouse pointer to (450, 600).
-    Python script:
-    import pyautogui
-    pyautogui.moveTo(450, 600)
-    pass
-def write(str=stringType, interval=secs_between_keys):
-    """writes the string wherever keyboard cursor is at the function calling time with (secs_between_keys) seconds between characters
-    Example:
-    High Level Goal: Write "Hello world" with 0.1 seconds rate
-    Python script:
-    import pyautogui
-    pyautogui.write("Hello world", 0.1)
-    pass
-def press(str=string_to_type):
-    """simulate pressing a key down and then releasing it up. Sample keys include ’enter’, ’shift’, arrow keys, ’f1’.
-    Example:
-    High Level Goal: Press enter key now
-    Python script:
-    import pyautogui
-    pyautogui.press("enter")
-    pass
-def hotkey(*args = list_of_hotkey):
-    """Keyboard hotkeys like Ctrl-S or Ctrl-Shift-1 can be done by passing a list of key names to hotkey(). Multiple keys can be pressed together with a hotkey.
-    Example:
-    High Level Goal: Use Ctrl and V to paste from clipboard
-    Python script:
-    import pyautogui
-    pyautogui.hotkey("ctrl", "v")
-    pass
-Here are a few examples for generating the script and format to be followed:
-Example 1:
-{RETREIVED_EXAMPLE_1}
-Example 2:
-{RETREIVED_EXAMPLE_2}
-Example 3:
-{RETREIVED_EXAMPLE_3}
-Example 4:
-{RETREIVED_EXAMPLE_4}
-Example 5:
-{RETREIVED_EXAMPLE_5}
-Here are the Textual Elements found on the screen along with their coordinates:
-{TEXTUAL_ELEMENTS_FROM_DETACT}
-Here are the Icon/Image Elements found on the screen along with their coordinates:
-{ICON_ELEMENTS_FROM_DETACT}
-Here are the Color Elements found on the screen along with their coordinates:
-{COLOR_ELEMENTS_FROM_DETACT}
-Based on the list of UI elements and the examples above, generate the pyautogui script for the following task:
-{TASK}
+Reliance on closed models like GPT-4V poses integration challenges due to high costs and time constraints.
+Despite efforts for equal representation and data collection without personal information, biases may be introduced as the dataset is exclusively in English, and human-curated content may have temporal biases.
+10 Ethics Statement
+As a part of the dataset creation process, we carefully review the pipeline at every stage, ensuring there is no personally identifiable information or offensive content, either during data collection or through the use of LLMs.
+For all purposes, we create dummy accounts that mimic real-like user content.
+To get the gold labels scripts we seek help from well-qualified student workers, approved through the institution, and get the bounding box data annotated through MTurk workers, both of whom are paid \$$\$25$ per hour, which is greater than the minimum wage rate (We detail this process in the supplementary material).
+Human studies are also done with the help of student workers approved by the institution at the above-mentioned payscale.
+We also ensure that all groups have equitable representation and that no personal opinions are reflected in the dataset, avoiding bias during the collection as well as the annotation process.
+Acknowledgements
+We extend our heartfelt gratitude to Writer.com for their generous and unwavering support throughout this project.
+Their dedicated team’s expertise and collaboration were invaluable in achieving our goals.